OpenStudy (anonymous):

A box has 1 red marble, 3 blue marbles, and 4 green marbles. Maya draws a blue marble randomly from the box, replaces it, and then draws another marble randomly. What is the probability of drawing 2 blue marbles in a row? Explain your answer.

OpenStudy (anonymous):

Since the marble is replaced between draws, each draw is independent. To find the probability of two independent events, you multiply the individual probabilityes together. Since P(blue) = 4/8 = 1/2 the probability of two blues in a row is (1/2) * (1/2) = 1/4